# Resilient Alerting Protocols for Blockchains

### v1

@@ Appendix D Simultaneous Alerting Game - $\mathsf{Alert}$ Dominance for Low Bribes
Claim 3 (restated). In the simultaneous alerting game, if the adversary’s offered bribe for a node $i$ is $\beta_{i}\leq{{\lambda}}$, then the action $\mathsf{Alert}$ weakly dominates $\mathsf{NoAlert}$ for node $i$. And $\mathsf{Alert}$ is strictly dominant if $\beta_{i}<{{\lambda}}$.
@@ Appendix E Simultanious Alerting Game - Mixed Strategy Equilibria
Multiplying Equation 12 by $q_{i}$ and using $q_{i}Q_{-i}=\prod_{j=1}^{n}q_{j}=Q$} yields $\beta_{i}q_{i}\geq{{\lambda}}\,(n-1)\,Q$ for all such $i$. Summing over $i=1,\ldots,n$, we get
@@ Appendix G TEE-based protocol - Alert Capability
Consider any node $i$. Local computation to form a commitment ${\gamma}_{i}$ is negligible compared to $\Delta_{\sf write}$. If node $i$ sends her transaction at any time $t\in[t_{0},\,t_{0}+\Delta_{\sf commit}-\Delta_{\sf write}]$}, then by the synchrony assumption the transaction is included in some block by time $t+\Delta_{\sf write}\leq t_{0}+\Delta_{\sf commit}$}, i.e., before the end of the commit window.
@@ Appendix I Deterministic Node Sequencing
The protocol uses a straightforward lexicographic next-permutation algorithm (Worlton, 1968) to sequence the nodes. Each node and the smart contract maintain the permutation state representing the current order of nodes. The sequence begins with the canonical permutation $\pi_{1}=(1,2,\ldots,n)$, corresponding to the lexicographically smallest ordering of all $n$ nodes. Both nodes and the contract implement the same algorithm: given the current permutation $\pi_{t}$, the function $\textit{next\_perm}({\cdot})$ computes $\pi_{t+1}=\textit{next\_perm}({\pi_{t}})$} by identifying the rightmost pair $(i,j)$ such that $\pi_{t}(i)<\pi_{t}(j)$, swapping these elements, and reversing the suffix following index $i$. This operation deterministically produces the next permutation in lexicographic order and requires no communication among nodes or with the smart contract. This ensures that all nodes and the contract have a consistent view of the current permutation.

### v2

@@ Appendix D Simultaneous Alerting Game - $\mathsf{Alert}$ Dominance for Low Bribes
Claim 3 (restated). In the simultaneous alerting game, if the adversary’s offered bribe for a node $i$ is $\beta_{i}\leq{{\lambda}}$, then the action $\mathsf{Alert}$ weakly dominates $\mathsf{NoAlert}$ for node ii. And $\mathsf{Alert}$ is strictly dominant if $\beta_{i}<{{\lambda}}$.
@@ Appendix E Simultanious Alerting Game - Mixed Strategy Equilibria
Multiplying Equation 12 by $q_{i}$ and using $q_{i}Q_{-i}=\prod_{j=1}^{n}q_{j}=Q$} yields $\beta_{i}q_{i}\geq{{\lambda}}\,(n-1)\,Q$ for all such ii. Summing over $i=1,\ldots,n$, we get
@@ Appendix G TEE-based protocol - Alert Capability
Consider any node ii. Local computation to form a commitment ${\gamma}_{i}$ is negligible compared to $\Delta_{\sf write}$. If node $i$ sends her transaction at any time $t\in[t_{0},\,t_{0}+\Delta_{\sf commit}-\Delta_{\sf write}]$}, then by the synchrony assumption the transaction is included in some block by time $t+\Delta_{\sf write}\leq t_{0}+\Delta_{\sf commit}$}, i.e., before the end of the commit window.
@@ Appendix I Deterministic Node Sequencing
The protocol uses a straightforward lexicographic next-permutation algorithm (Worlton, 1968) to sequence the nodes. Each node and the smart contract maintain the permutation state representing the current order of nodes. The sequence begins with the canonical permutation $\pi_{1}=(1,2,\ldots,n)$, corresponding to the lexicographically smallest ordering of all $n$ nodes. Both nodes and the contract implement the same algorithm: given the current permutation $\pi_{t}$, the function $\textit{next\_perm}({\cdot})$ computes $\pi_{t+1}=\textit{next\_perm}({\pi_{t}})$} by identifying the rightmost pair $(i,j)$ such that $\pi_{t}(i)<\pi_{t}(j)$, swapping these elements, and reversing the suffix following index ii. This operation deterministically produces the next permutation in lexicographic order and requires no communication among nodes or with the smart contract. This ensures that all nodes and the contract have a consistent view of the current permutation.